OpenStudy (anonymous):

if a and b are integers and 14/3=a/b which could be the value of a+b

OpenStudy (anonymous):

Note that \(\dfrac{14}{3}=\dfrac{-14}{-3}\). So if this is equal to \(\dfrac{a}{b}\), then \(a=14\) and \(b=3\), or \(a=-14\) and \(b=-3\).

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

what a happy face